OpenStudy (loser66):

Use Implicit function theorem to show F(x,y) = y^2+2xy +x-1 =0 can be solved for y in terms of x near all point (x0,y0) with \(y_0\neq -x_0\) Check the result by a) directly (using algebra) b) determine if F(x,y)=0 is locally solvable near any point \((x_0,-x_0). Please, help

OpenStudy (loser66):

IFT says that we can solve for y in terms of x if \(\dfrac{\partial F}{\partial y}\neq 0\), and we have it is \(2y +2x\), hence if \(y_0\neq -x_0\), it is not equal 0, hence we are done.

OpenStudy (loser66):

For part a) I am not sure what I am supposed to do. If checking directly by using algebra, I have \(y = -2x\pm\sqrt{x^2-x+1}\) But I don't know how to argue more.

OpenStudy (loser66):

@tkhunny

OpenStudy (tkhunny):

What more is there for Part A? Looks to me like you're done.

OpenStudy (loser66):

I have 2 different equations, right? \(\pm \sqrt....\) how to argue?

OpenStudy (loser66):

And how to do b?

OpenStudy (tkhunny):

"\(\huge{can}\) be solved for y in terms of x" Throw one out. Is x^2 - x + 1 < 0 anywhere?

OpenStudy (loser66):

no solution for that

OpenStudy (tkhunny):

Then, it always exists in Real numbers and we are done with the existence and our ability to solve this one. Substitute for \(y_{0}\) as indicated.

OpenStudy (loser66):

Thank you so much.

OpenStudy (tkhunny):

Once in a while, I get one. :-)